OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7.

OpenStudy (anonymous):

y = 4x + 10 y = one fourthx + 5 halves y = -4x - 6 y = -one fourthx + 3 halves

OpenStudy (danjs):

y = (slope)*x + (intercept) y = m *x + b

OpenStudy (danjs):

If one line is parallel to another, they will have the same slope m

OpenStudy (anonymous):

okay

OpenStudy (danjs):

so the slope is 4 and they gave you a point (x,y) = (-2 , 2) You know everything but the y intercept now, b y = m*x + b plug in and solve

OpenStudy (danjs):

2 = 4*(-2) + b

OpenStudy (anonymous):

2=-8+b

OpenStudy (danjs):

yes, so the line parallel has slope --- m = 4 y-intercept --- b = 10 y = m*x + b

OpenStudy (anonymous):

y = 4x + 10 y = one fourthx + 5 halves y = -4x - 6 y = -one fourthx + 3 halves

OpenStudy (anonymous):

so a

OpenStudy (danjs):

yep

OpenStudy (anonymous):

okay awesome thanks

OpenStudy (danjs):

no prob.